OpenStudy (anonymous):

Let u = <-7, -2>. Find 4u.

ganeshie8 (ganeshie8):

simply multiply each component of by 4

OpenStudy (anonymous):

i got <28, -8> im not sure if that is right

ganeshie8 (ganeshie8):

hmm what happened to the \(-\) sign of first component ?

OpenStudy (anonymous):

oh! so it will be -28, -8?

ganeshie8 (ganeshie8):

<-28, -8> is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thank! can you help me with a few other questions i got wrong on my homework?

ganeshie8 (ganeshie8):

I'll try, ask...

OpenStudy (anonymous):

Given that P = (5, 9) and Q = (13, 12), find the component form and magnitude of ->over PQ

OpenStudy (anonymous):

i had problems with this and just randomly answered <-8, -3>, 73

ganeshie8 (ganeshie8):

\(\large \overrightarrow{PQ} = \overrightarrow{OQ} - \overrightarrow{OP} = \langle 13, 12\rangle - \langle 5, 9\rangle = ?\)

OpenStudy (anonymous):

i am confused?

ganeshie8 (ganeshie8):

to subtract two vectors, we just subtract each components : \(\large \overrightarrow{PQ} = \overrightarrow{OQ} - \overrightarrow{OP} = \langle 13, 12\rangle - \langle 5, 9\rangle = \langle 13-5, ~12-9 \rangle = ? \)

ganeshie8 (ganeshie8):

All we're doing is subtraction only :) 13 - 5 = ? 12 - 9 = ?

OpenStudy (anonymous):

8 and 3?

OpenStudy (anonymous):

but i am looking at the answer choices and they include 73 or sqrt 73

ganeshie8 (ganeshie8):

Yes ! \(\large \overrightarrow{PQ} = \langle 8, ~3 \rangle \)

ganeshie8 (ganeshie8):

thats the component form of PQ ^

ganeshie8 (ganeshie8):

You need to find its magnitude still

ganeshie8 (ganeshie8):

magnitude = \(\large \sqrt{8^2 + 3^2} = ?\)

OpenStudy (anonymous):

14?

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

magnitude = \(\large \sqrt{8^2 + 3^2} = \sqrt{64 + 9} = \sqrt{73}\)

OpenStudy (anonymous):

so <8, 3>, \[\sqrt{73}\]

OpenStudy (anonymous):

what about Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v, I answered <-36, -18>

OpenStudy (anonymous):

this one i am very confused by

ganeshie8 (ganeshie8):

multiply u by 9 multiply v by 6 subtract

ganeshie8 (ganeshie8):

u = <-5, 1> 9u = ?

OpenStudy (anonymous):

<-5, 1> i subtract that?

ganeshie8 (ganeshie8):

multiply it by 9

ganeshie8 (ganeshie8):

u = <-5, 1> 9u = 9*<-5,1> = <-45, 9>

ganeshie8 (ganeshie8):

see if u can find `6v`

OpenStudy (anonymous):

<63, -36>?

ganeshie8 (ganeshie8):

Nope. v = <7, -4> 6v = 6*<7, -4> = ?

OpenStudy (anonymous):

<42, -24>?

ganeshie8 (ganeshie8):

Yes ! now subtract

ganeshie8 (ganeshie8):

9u - 6v = <-45, 9> - <42, -24> = ?

OpenStudy (anonymous):

so -54, 66?

OpenStudy (anonymous):

this is where i get confused a little because i subtract the neg 45 from 9? or do i subtract the neg 54 and 66?

ganeshie8 (ganeshie8):

you need to subtract the corresponding components

ganeshie8 (ganeshie8):

9u - 6v = <\(\color{red}{-45}\), \(\color{green}{9}\)> - <\(\color{red}{42}\), \(\color{green}{-24}\)>

OpenStudy (anonymous):

<-87, 33> ???

ganeshie8 (ganeshie8):

9u - 6v = <\(\color{red}{-45}\), \(\color{green}{9}\)> - <\(\color{red}{42}\), \(\color{green}{-24}\)> = <\(\color{red}{-45-42}\), \(\color{green}{9--24}\)> = <\(\color{red}{-87}\), \(\color{green}{33}\)>

ganeshie8 (ganeshie8):

Yes ! looks you nailed it !!

OpenStudy (anonymous):

thank you! im going to try Let u = <9, 4>, v = <-2, 5>. Find u + v. can you tell me if i get the answer right?

OpenStudy (anonymous):

okay i got <7,9>? @ganeshie8

ganeshie8 (ganeshie8):

Excellent !!

ganeshie8 (ganeshie8):

you've become the maestro of vectors !!

OpenStudy (anonymous):

yay, thank you very much!

ganeshie8 (ganeshie8):

np, you're welcome :)